OpenStudy (anonymous):

Find the equation of the axis of symmetry: f(x)=3(x-9)^2+2 Answers: A. x=9 B. x=-2 C. x=2 D. x=-9 and E. None of the Above

jimthompson5910 (jim_thompson5910):

Hint: The axis of symmetry of \[\Large y = a(x-h)^2 + k\] is \[\Large x = h\]

OpenStudy (anonymous):

i don't know what to do with that hint?

jimthompson5910 (jim_thompson5910):

Example: The equation \[\Large y = 5(x-2)^2 + 14\] is in the form \[\Large y = a(x-h)^2 + k\] So we see that a = 5, h = 2 and k = 14 So the axis of symmetry of \[\Large y = 5(x-2)^2 + 14\] is \[\Large x = 2\]

OpenStudy (anonymous):

okay i got it now. Thank you.

jimthompson5910 (jim_thompson5910):

you're welcome